OpenStudy (anonymous):

Let y = f(x) be a function with domain D = [−2, 6] and range R = [−4, 8]. Find the domain D and range R for each function. Assume f(2) = 8 and f(6) = −4. (Enter your answers using interval notation.)

OpenStudy (anonymous):

a) y = −2f(x) domain D = range R = (b) y = f 1 2 x domain D = range R = (c) y = f(x − 3) + 1 domain D = range R = (d) y = f(x + 2) − 3 domain D = range R = (e) y = f(−x) domain D = range R = (f) y = −f(x) domain D = range R = (g) y = |f(x)| domain D = range R =